DECELERATION AND TRAPPING OF HEAVY DIATOMIC MOLECULES FOR PRECISION MEASURE-MENTS

J. E. VAN DEN BERG, S. N. HOEKMAN TURKESTEEN, E. B. PRINSEN, \underline{S. HOEKSTRA}, Zernikelaan 25, 9747 AA, Groningen, The Netherlands.

We are setting up a novel type of Stark-decelerator optimized for the deceleration and trapping of heavy diatomic molecules. Aim of these experiments is to prepare a trapped sample of ultracold molecules for precision studies of fundamental symmetries. The decelerator uses ring-shaped electrodes to create a moving trapping potential, a prototype of which has been shown to work for CO molecules^{a}. Molecules can be decelerated and trapped in the weak-field seeking part of excited rotational states. The alkaline-earth monohalide molecules (currently we focus on the SrF molecule) are prime candidates for next generation parity violation and electron-EDM studies^{b}. We plan to combine the Stark deceleration with molecular laser cooling to create a trapped sample of molecules at a final temperature of ∼ 200 μK.

aA. Osterwalder, S. A. Meek, G. Hammer, H. Haak and G. Meijer Phys. Rev. A \underline{81} (51401), 2010.

bT. A. Isaev, S. Hoekstra, R. Berger Phys. Rev. A \underline{82} (52521), 2010